OpenStudy (jenniferjuice):

IM TIRED AND I NEED HELP SO SOMEONE, ANYONE HELP ME! ANYONE WELCOME! JUST COME AND HELP! MEDALS AWARDED !

OpenStudy (jenniferjuice):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

OpenStudy (anonymous):

the only one that fulfills y=2 when x=3 and y=6 when x=-9 is the second one

OpenStudy (texaschic101):

find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (3,2) -- x1 = 3and y1 = 2 (-9,6) -- x2 = -9 and y2 = 6 now plug in your info into the slope formula and find your slope

OpenStudy (anonymous):

it should be y = -1/3x + 3

OpenStudy (anonymous):

which would be equation #2

OpenStudy (jenniferjuice):

oh okay thanks ! :D all of you's haha

OpenStudy (jenniferjuice):

anyone wanna stick around and help me with a couple more ..hehe